What is the mass of H2O that will be produced from 2.5 moles of O2?
1
Expert's answer
2020-09-30T07:45:36-0400

At first, the balance equation should be written:

O2+2H2=2H2O.

It is seen that number of moles of water is 2 times higher, so n=2×2.5=5 moles of water.

The mass of water may be found from the formula:

m=n×M,

M=2×1+16=18(g/mol),

m=5×18=90(g).